OpenStudy (anonymous):

Which shows the correct word and standard forms for the decimal? 20 + 6 + 0.09 A. twenty-six and nine tenths 26.9 B. twenty-six and nine hundredths 26.09 C. twenty-six and nine hundredths 26.9 D. twenty-six and nine thousandths 26.009

OpenStudy (jbo11):

well 1/10 = .1 and 1/100 = .01 so .09 = 9/100

OpenStudy (anonymous):

ok i get it now thx!

OpenStudy (jbo11):

np